# MAC: A Benchmark for Multiple Attribute Compositional Zero-Shot Learning

### v2

@@ 3.3 Task Formulation
Given the attribute set $A=\{a_{1},a_{2},\cdots,a_{|A|}\}$ and object set $O=\{o_{1},o_{2},\cdots,o_{|O|}\}$ as the semantic primitives, where ||$\cdot$|| ⋅ | denotes the number of elements in the set, the multi-attribute compositional label space is defined as $C=\left\{\left\langle S,o\right\rangle\mid S\subseteq A,o\in O\right\}$. We define the set of seen compositions and unseen compositions as $C^{s}\subseteq C$ and $C^{u}\subseteq C$, respectively, where $C^{s}\cap C^{u}=\emptyset$. During training, only the data of seen compositions is provided. Implementing Image-to-Text retrieval for multi-attribute compositions is impractical due to the massive search space. Therefore, we design a multi-label single attribute composition classification task for our dataset. Notably, multiple labels apply only to the attribute part. This task aims to predict compositions in the test space $C^{t}$ in both closed-world setting and open-world setting. In closed-world setting, the solution space is defined as $C^{t}=\left\{\langle a,o\rangle\mid\exists\langle S,o\rangle\in C^{s}\cup C^{u},\,a\in S\right\}$. Additionally, we focus more on the open-world setting because it is more consistent with real-world scenarios. In this setting, the solution space is defined as the Cartesian product of the attribute set and the object set, i.e., $C^{t}=|A|\times|O|$.
3.4 Evaluation Metric
@@
To evaluate the accuracy and comprehensiveness of model respectively, we use Exact Match, Top-1 precision (Top1-P), Top-5 Recall (Top5-R) and Coverage following  [47, 52]. Exact Match is the proportion of instances that all true labels are ranked higher than any false labels. The Top1-P reflects the model’s ability to correctly predict compositions. The Top5-R evaluates the ability of a model to identify relevant labels within its top 5 predictions. Coverage quantifies the model’s ability to capture all true labels by computing the smallest integer $K$ such that the top $K$ predicitons cover all true labels of an instance. We also introduce Top-1 Precision for attributes (Top1-P-attr) and objects (Top1-P-obj) to measure the model’s performance on individual primitives. Calculation details for these metrics are in section 1.3 of the Appendix . To ensure consistency with prior work, we also report results on AUC, Seen, and Unseen following  [30].
4 Proposed Method
@@
Figure 6: Overview of the proposed method. MVP-Integrator is the first CZSL method that abandons composition input and automatically learn the associations between primitives. $\mathcal{L}_{a}$ is the BCE loss on attribute classification and $\mathcal{L}_{o}$ is Cross-Entropy loss for object classification.
@@
Compared to previous benchmarks, multi-attribute compositional learning is more challenging as it requires the model to identify all feasible attributes. To achieve accurate composition predictions, a model must disentangle semantic primitives to independently learn attributes and objects from composition data. It should also effectively capture relationships between primitives, including attribute-attribute and attribute-object associations. However, existing models struggle with multi-attribute composition learning since they are designed for single-attribute datasets, ignoring the relationship between different attributes. We propose Multi-attreibute Visual-Primitive Integrator (MVP-Integrator) to adapt CLIP [34] for multi-attribute composition learning. MVP-Integrator is the first CZSL model to eliminate compositional inputs and instead uses dual-branch prompt tuning for primitive disentanglement and leverages self-attention to capture relationships between primitives automatically.
@@
Prompt Tuning. Unlike previous methods, we discard compositional text inputs and instead employ a dual-branch prompt tuning strategy with independent learnable prompts for attributes and objects. For an attribute-object composition $(a\;o)$, we convert the natural language prompt “a photo of ${a}$ ${o}$” into two independent prompt: $P_{a}=[p^{a}_{1},p^{a}_{2},\dots,p^{a}_{K},v_{a}]$ and $P_{o}=[p^{o}_{1},p^{o}_{2},\dots,p^{o}_{K},v_{o}]$ , where $[p_{1},p_{2},\dots,p_{K}]$ are the trainable embedding and $v_{a},v_{o}$ are the frozen word vocabulary embeddings of $a,o$.
Multi-Attribute Visual-Primitive Association. Due to the complexity of the compositional space (the cartesian product of the attribute and object sets), existing methods often require specialized modules for attribute-object association and separate mechanisms for text-image feature fusion.
@@
In contrast, our method eliminate the compositional branch significantly reducing the complexity of text features. This enables us to employ a lightweight transformer encoder to jointly model relationships between attributes and objects, inter-attribute correlations, and cross-modal interactions between vision and language. Specifically, let $I$ denote the concatenated image patch representations and the class token representation extracted by the CLIP’s visual encoder. $T$ represents the concatenated text representations extracted by CLIP’s text encoder for all primitives. The combined tokens $M=[I;T]$ are fed into a lightweight transformer encoder to model both the relationship between different semantic primitives (attributes and objects) and the relationship between the image and the text. Finally, the cosine similarity between refined image class token representation and the original text token representations $T$ is computed. It’s worth noting that the methods without a composition branch do not require training on all possible compositions. It significantly reduces the computational resources for training and inference on the combined dataset: the input number to the text encoder drops from $|A|\times|O|$ to $|A|+|O|$.
5 Experiments
@@
Table 3: Main results on MAC. CGE [30] uses ResNet18 pretrained on ImageNet, while other methods use CLIP ViT-L/14. Results of LLAVA [23] is obtained by probabilistic inference techniques from [55]. †  indicates that the results are obtained without training on MAC.
5.1 Implementation Details
We evaluate methods on the multi-label single composition task described in Sec. 3.3 with metrics in Sec. 3.4. CGE uses a ResNet18 pretrained on ImageNet, while other methods use a CLIP ViT-L/14 backbone. Troika is trained and evaluated on an NVIDIA L20 GPU, whereas other methods use an RTX 4090 GPU. Hyperparameter details are provided in section 2 of the Appendix. All the methods are trained and evaluated in the same way. During training, we use cross-entropy loss for the object branch and BCE loss for the multi-label branch (attribute branch and composition branch). During inference, we add all the predictions to get the final composition prediction:
@@ 5.1 Implementation Details
For methods that lack one or more branches, we disregard the missing branches during both training and inference. We also evaluate visual-language models (VLMs) like LLaVA [23]. We use a probabilistic inference technique to fully leverage their potential. We calculate class probabilities by summing the token probabilities for each class name and rank prob(class name|image, prompt) as the final prediction [5, 55]. Directly calculating these probabilities in the composition space is infeasible. Instead, we compute the probabilities for all object and attribute categories for each image and sum them to derive the composition probability.
5.2 Main Results
We experiment with seven models on MAC. CGE [30] is the only one not based on CLIP. CLIP [34], CoOP [58], CSP [31], DFSP [24], GIPCOL [49], and Troika [11] are all common CZSL baselines based on the CLIP ViT-L/14 backbone. LLaVA [23] is the most widely used VLM, introducing more parameters compared to other models.
Tab. 3 shows the performance of various models in both the closed-world and open-world settings on the MAC dataset. In the closed-world setting, MVP-Integrator significantly outperforms other methods, achieving 17.42% Exact Match, 63.45% Top1-P, 62.37% Top5-R and 27.89 on AUC. In the open-world setting, our method also achieves the best results in all metrics. Notably, MVP-Integrator maintains consistent performance across both settings, while other methods experience a significant drop in performance in the open-world setting. Meanwhile, the performance of CSP is even worse than the untrained CLIP. LLaVA has high prediction accuracy but performs poorly in overall prediction quality, as measured by the Top5-R and Coverage.
Compared with the closed-world setting, the open-world setting lacks infeasible composition priors, making it more challenging and better aligned with real-world applications. Methods based on CLIP show a significant drop in performance from closed-world to open-world settings, with a relative 22.01% drop in Exact Match. Specifically, the performance of models based on CLIP in terms of the Top1-P-attr drops by at least 3.98%, while the drop in terms of the Top1-P-obj is just 0.22%. This indicates that existing methods are heavily influenced by composition priors and have a weaker understanding of the essential characteristics of attributes, resulting in poor open-world capabilities. In contrast, our MVP-Integrator shows a difference of only relatively 3.7% in Exact Match and 0.63% in Top1-P-attr between different settings, with overall results remaining nearly unchanged. This demonstrates that decoupling the learning of attribute and object knowledge and modeling the correlations between multiple attributes are crucial for significantly improving the model’s open-world abilities.
5.3 Efficiency analysis
@@
Table 4: Efficiency analysis on open-world setting. FLOPs quantify the computational load for a model on a single image. GPU memory is measured with a batch size of 16. Inference time is averaged over 1k randomly selected MAC samples.
As shown in Tab. 4, our MVP-Integrator achieves 25$\times$ faster inference speed while using 15% GPU memory compared to existing methods in open-world settings. This improvement comes from addressing a key limitation of current approaches: they require checking every possible compositions of attributes $|A|$ and object categories $|O|$, leading to a solution space of $|A|\times|O|$ compositions. This results in significant GPU memory usage and slower inference speed as they need to extract all text features for every possible composition and execute cross-modal fusion. In contrast, the MVP-Integrator eliminate the composition branch. This makes it more flexible, with a linear solution space of $|A|+|O|$.
5.4 Cross-dataset Evaluation
@@
Table 5: Cross-Dataset Evaluation Results. Models trained on MAC are directly evaluated in a zero-shot manner on the test set.
Figure 7: Qualitative results on MIT-States [15]. We directly test the MAC-trained MVP-Integrator on MIT-States data, finding that its top-3 attribute predictions introduce suitable new attributes while retaining the original ground truth.
@@
Tab. 5 shows the results of cross-dataset evaluation. We can observe that the models trained in MAC achieve comparable results with their counterparts trained on MIT-States. The cross-dataset performance of MVP-Integrator and Tri-CoOP even outperforms their counterparts. This indicates that the multiple comprehensive attribute labels of MAC help to improve the quality of the model’s composition predictions. In Fig. 7, we use MVP-Integrator trained on the MAC to generate the top three predictions on images from the MIT-States. MVP-Integrator provides representative and relevant new attributes while preserving the original annotations in its top predictions. These newly identified attributes provide a more detailed description of appearance compared to the single-attribute annotation in MIT-States. Furthermore, although the model’s predictions are reasonable, they are often misjudged as incorrect due to the single-attribute annotation of MIT-States. This underscores how single attribute annotations can lead to inaccurate evaluation.
5.5 Ablation Studies
In Tab. 6, we conduct extensive experiments and report the closed-world results to show the effectiveness of our design.
@@ 5.5 Ablation Studies
Table 6: Ablation study on MVP-Integrator. Group 1 is the zero-shot result of CLIP. “Single” represents the composition branch, and “Double” indicates the branches for attribute and object. “CLS” and “Patch” refer to the source of image token fed into the multimodal fusion module. Group 7 is the full version of MVP-Integrator.
Disentangling Primitives. Group 3 outperforms Group 2 significantly in Exact Match (9.20% vs. 6.53%). This improvement underscores the benefit of disentangling primitives, showing that predicting attributes and objects separately is more effective than predicting compositions.
@@
Multi-attribute Visual-Primitive Association. The integration of the Cross-Modal Fusion module results in a obvious improvement in Exact Match compared to the models without this module (Group 4 vs Group 3). This highlights the importance of modeling the relationships across different primitives and interactions between images and language.
@@
CLS Token vs. Patch Token. Group 5 performs worse than Group 4, especially in terms of attribute accuracy. This indicates that the CLS token of the CLIP image encoder plays a more significant role in enhancing the image features, as it has been directly trained to align with text features.
@@
Patch Token Contributions. Group 7 outperforms Group 4 in all three metrics. This indicates that Patch Token features contribute valuable information, enhancing the model’s ability to understand semantic primitives.
6 Discussion and Conclusion

### v4

@@ III-D Task Formulation
Given the attribute set $A=\{a_{1},a_{2},\cdots,a_{|A|}\}$ and object set $O=\{o_{1},o_{2},\cdots,o_{|O|}\}$ as the semantic primitives, where $|\cdot|$ denotes the number of elements in the set, the multi-attribute compositional label space is defined as $C=\left\{\left\langle S,o\right\rangle\mid S\subseteq A,o\in O\right\}$. We define the set of seen compositions and unseen compositions as $C^{s}\subseteq C$ and $C^{u}\subseteq C$, respectively, where $C^{s}\cap C^{u}=\emptyset$. During training, only the data of seen compositions is provided. Implementing Image-to-Text retrieval for multi-attribute compositions is impractical due to the massive search space. Therefore, we design a multi-label single attribute composition classification task for our dataset. This task aims to predict compositions in the test space $C^{t}$ in both closed-world setting and open-world setting. In closed-world setting, the solution space is defined as $C^{t}=\left\{\langle a,o\rangle\mid\exists\langle S,o\rangle\in C^{s}\cup C^{u},\,a\in S\right\}$. Additionally, we focus more on the open-world setting because it is more consistent with real-world scenarios. In this setting, the solution space is defined as the Cartesian product of the attribute set and the object set, i.e., $C^{t}=A\times O$.
III-E Evaluation Metric
@@
To evaluate the accuracy and comprehensiveness of model respectively, we use Exact Match, Top-1 precision (Top1-P), Top-5 Recall (Top5-R) and Coverage following  [44, 45]. Exact Match is the proportion of instances that all true labels are ranked higher than any false labels. The Top1-P reflects the model’s ability to correctly predict compositions. The Top5-R evaluates the ability of a model to identify relevant labels within its top 5 predictions. Coverage quantifies the model’s ability to capture all true labels by computing the smallest integer $K$ such that the top $K$ predicitons cover all true labels of an instance. We also introduce Top-1 Precision for attributes (Top1-P-attr) and objects (Top1-P-obj) to measure the model’s performance on individual primitives. To ensure consistency with prior work, we also report results on AUC, Seen, and Unseen following  [11]. We use the top1-P to compute seen and unseen accuracies.
IV Proposed Method
@@
Figure 7: The proposed method. $\mathcal{L}_{a}$ is BCE loss on attribute classification and $\mathcal{L}_{o}$ is Cross-Entropy loss for object classification.
@@
Compared to previous benchmarks, multi-attribute compositional learning is more challenging as it requires the model to identify all feasible attributes. To achieve accurate composition predictions, a model must disentangle semantic primitives to independently learn attributes and objects from composition data. It should also effectively capture relationships between primitives, including attribute-attribute and attribute-object associations. However, existing models struggle with multi-attribute composition learning since they are designed for single-attribute datasets, ignoring the relationship between different attributes. We propose Multi-attreibute Visual-Primitive Integrator (MVP-Integrator) to adapt CLIP [37] for multi-attribute composition learning, as shown in Fig.˜7. MVP-Integrator is the first CLIP-based CZSL model to eliminate compositional inputs and instead uses dual-branch prompt tuning for primitive disentanglement and leverages self-attention to capture relationships between primitives automatically.
@@
Prompt Tuning. Unlike previous methods, we discard compositional text inputs and instead employ a dual-branch prompt tuning strategy with independent learnable prompts for attributes and objects. For an attribute-object composition (ao)(a\;o), we convert the natural language prompt “a photo of a{a} o{o}” into two independent prompt: $P_{a}=[p^{a}_{1}\;p^{a}_{2}\;\dots\;p^{a}_{K}\;v_{a}]$ and $P_{o}=[p^{o}_{1}\;p^{o}_{2}\;\dots\;p^{o}_{K}\;v_{o}]$ , where $p_{1}\;p_{2}\;\dots\;p_{K}]$ are the trainable embedding and $v_{a},v_{o}$ are the frozen word vocabulary embeddings of $a,o$. The trainable prompts are fed into the text encoder $\mathcal{F}_{text}$ to obtain the text representation of all primitives, formulated as
$t_{a_{i}}$ $=\mathcal{F}_{\text{text}}(P_{a_{i}}),\quad t_{o_{j}}=\mathcal{F}_{\text{text}}(P_{o_{j}}),$ (1)
$t_{A}$ $=[t_{a_{1}}\;\ldots\;t_{a_{|A|}}],\quad t_{O}=[t_{o_{1}}\;\ldots\;t_{o_{|O|}}].$
@@
Multi-Attribute Visual-Primitive Association. Due to the complexity of the compositional space (the cartesian product of the attribute and object sets), existing methods often require specialized modules for attribute-object association and separate mechanisms for text-image feature fusion. In contrast, our method eliminate the compositional branch significantly reducing the complexity of text features. This enables us to employ a lightweight transformer encoder to jointly model relationships between attributes and objects, inter-attribute correlations, and cross-modal interactions between vision and language. Specifically, let $I$ denote the concatenated image patch representations and $C$ denote the class token representation extracted by the CLIP’s visual encoder. $T=[t_{A};t_{O}]$ represents the concatenated text representations extracted by CLIP’s text encoder for all attributes and objects. The combined tokens $M=[C;I;T]$ are fed into a lightweight transformer encoder to model both the relationship between different semantic primitives (attributes and objects) and the relationship between the image and the text. Finally, the cosine similarity is calculated between the refined image class token $C^{\prime}$ and the original text token representations $T$. It’s worth noting that the methods without a composition branch do not require training on all possible compositions. It significantly reduces the computational resources for training and inference on the combined dataset: the input number to the text encoder drops from $|A|\times|O|$ to $|A|+|O|$.
Training objective. As shown in Fig.˜7, MVP-Integrator contains two branches for attributes and objects respectively. Given the text representation for all primitives and visual features, the logits of of assigning labels of the attribute $a_{i}$, object $o_{j}$ can be computed separately as
[…]
The attribute branch is trained with BCE loss, formulated as:
$\mathcal{L}_{a}=-\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\sum_{i=1}^{|A|}\big[$ $y_{xi}\log\sigma(s_{x,a_{i}})$ (3)
$+(1-y_{xi})\log(1-\sigma(s_{x,a_{i}}))\big],$
where $y_{xi}\in{0,1}$ indicates the presence of attribute $a_{i}$ of image $x$, and $\sigma(\cdot)$ is the sigmoid activation function. The object branch is trained using the standard Cross-Entropy (CE) loss:
$\mathcal{L}_{o}=-\frac{1}{|\mathcal{X}|}\sum_{x\in\mathcal{X}}\log\left(\frac{\exp(s_{x,o_{y_{x}}})}{\sum_{j=1}^{|O|}\exp(s_{x,o_{j}})}\right),$ (4)
where $o_{y_{x}}$ denotes the labeled object class for image $x$. Therefore, the overall training loss $\mathcal{L}$ is defined as:
$\mathcal{L}=\mathcal{L}_{a}+\mathcal{L}_{o}.$ (5)
V Experiments
TABLE III: Main results on MAC. CGE [11] uses ResNet18 pretrained on ImageNet, while other methods use CLIP ViT-L/14. †  indicates that the results are obtained without training on MAC.
@@ V Experiments
V-A Implementation Details
We evaluate methods on the multi-label single composition task described in Section˜III-D with metrics in Section˜III-E. CGE uses a ResNet18 pretrained on ImageNet, while other methods use a CLIP ViT-L/14 backbone. Troika is trained and evaluated on an NVIDIA L20 GPU, whereas other methods use an RTX 4090 GPU. Hyperparameter details are provided in section D of Appendix. All the methods are trained and evaluated in the same way. During training, we use cross-entropy loss for the object branch and BCE loss for the multi-label branch (attribute branch and composition branch). During inference, we add all the predictions to get the final composition prediction:
@@ V-A Implementation Details
These probability distributions are derived by applying the softmax function to the respective image-text similarity scores. For methods that lack one or more branches, we disregard the missing branches during both training and inference. We do not use any special post processing (e.g., feasibility masking [2]) for open-world setting to directly measure the model’s compositional learning ability. We also evaluate visual-language models (VLMs) like LLaVA [16]. We use a probabilistic inference technique to fully utilize their potential. We calculate class probabilities by summing the token probabilities for each class name and rank prob(class name|image, prompt) as the final prediction [46, 47]. Calculating these probabilities directly in the composition space is infeasible. Instead, we compute the probabilities for all the classes of objects and attributes for each image and sum them to derive the composition probability.
V-B Main Results
We experiment seven models on MAC. CGE [11] is the only one not based on CLIP. CLIP [37], CoOP [12], CSP [13], DFSP [15], GIPCOL [34], and Troika [14] are all common CZSL baselines based on the CLIP ViT-L/14 backbone. LLaVA [16] is the most widely used VLM, introducing more parameters compared to other models.
Tab.˜III shows the performance of various models in both the closed-world and open-world settings on the MAC dataset. In the closed-world setting, MVP-Integrator significantly outperforms other methods, achieving 17.42% Exact Match, 63.45% Top1-P, 62.37% Top5-R and 27.89 on AUC. In the open-world setting, our method also achieves the best results. Notably, MVP-Integrator maintains consistent performance across both settings, while other methods show a significant drop in the open-world setting. LLaVA has high prediction accuracy but performs poorly in overall prediction quality, as measured by the Top5-R and Coverage. Compared with the closed-world setting, the open-world setting lacks infeasible composition priors, making it more challenging and better aligned with real-world applications. Methods based on composition branch show a significant drop in performance from closed-world to open-world settings, with a relative 22.01% drop in Exact Match. Specifically, the performance of models based on CLIP in terms of the Top1-P-attr drops by at least 3.98%. This indicates that existing methods are heavily influenced by composition priors and have a weaker understanding of the essential characteristics of attributes. In contrast, our MVP-Integrator shows a difference of only relatively 3.7% in Exact Match and 0.63% in Top1-P-attr between different settings, with overall results remaining nearly unchanged.
V-C Efficiency analysis
TABLE IV: Efficiency analysis on open-world setting. FLOPs quantify the computational load for a model on a single image. GPU memory is measured with a batch size of 16. Inference time is averaged over 1k randomly selected MAC samples.
@@ V-C Efficiency analysis
Figure 8: Qualitative results for cross-dataset evaluation. We directly test the MAC-trained MVP-Integrator on MIT-States [9] data, finding that its top-3 attribute predictions introduce suitable new attributes while retaining the original ground truth.
[…]
V-D Cross-dataset Evaluation
TABLE V: Cross-Dataset Evaluation Results. Models trained on MAC are directly evaluated on the test set.
@@ V-D Cross-dataset Evaluation
Tab.˜V shows the results of cross-dataset evaluation. We can observe that the models trained in MAC achieve comparable results with their counterparts trained on MIT-States, except for Troika. This discrepancy likely stems from Troika’s use of Adapter [48] to finetune CLIP’s visual encoder, which may have reduced its generalization ability. The cross-dataset performance of MVP-Integrator and Tri-CoOP even outperforms their counterparts. This indicates that the multiple comprehensive attribute labels of MAC help to improve the quality of the model’s composition predictions. In Fig.˜8, we use MVP-Integrator trained on the MAC to generate the top three predictions on images from the MIT-States. MVP-Integrator provides representative and relevant new attributes while preserving the original annotations in its top predictions. These newly identified attributes provide a more detailed description compared to the single-attribute annotation in MIT-States. Furthermore, although the model’s predictions are reasonable, they are often misjudged as incorrect due to the single-attribute annotation of MIT-States. This underscores how single attribute annotations can lead to inaccurate evaluation.
V-E Ablation Studies
In Tab.˜VI, we conduct extensive experiments and report the closed-world results to show the effectiveness of our design.
TABLE VI: Ablation study on MVP-Integrator. Group 1 is the zero-shot result of CLIP. “Single” represents the composition branch, and “Double” indicates the branches for attribute and object. “CLS” and “Patch” refer to the source of image token fed into the multimodal fusion module. Group 7 is the full version of MVP-Integrator.
@@ V-E Ablation Studies
Disentangling Primitives. Group 3, which uses separate prompt tuning branches for attributes and objects, significantly outperforms Group 2, which uses a single composition branch. This improvement underscores the benefit of disentangling primitives, showing that predicting attributes and objects separately is more effective than predicting compositions. To further support this finding, we conducted an additional experiment. When we added a composition branch (trained similarly to CoOp [1]) to our full MVP-Integrator, the performance dropped considerably, with the Exact Match score falling from 17.42% to 9.26% and the AUC from 27.89% to 24.65%. This reinforces our conclusion that a dual-branch structure is superior for this task and suggests that effectively leveraging a composition branch for multi-attribute compositional zero-shot learning remains an valuable research question.
@@
Multi-attribute Visual-Primitive Association. The integration of the Cross-Modal Fusion module results in an obvious improvement in Exact Match compared to the models without it (Group 4 vs. Group 3). This highlights the importance of modeling the interactions between images and language and relationships across different primitives. We further analyze the effect of primitive relationships in  Section˜V-F.
@@
CLS Token vs. Patch Token. To understand the source of visual information for our Integrator module, we compare two variants. Group 4 uses the single CLS token from CLIP’s image encoder. Group 5 uses the patch token and average the Integrator’s output tokens for the final prediction. Group 5 performs worse than Group 4, especially in terms of attribute accuracy. This indicates that the CLS token plays a more significant role in enhancing the image features, as it has been directly trained to align with text features.
@@
Patch Token Contributions. While the [CLS] token is crucial, local features from patch tokens also provide valuable information. Group 7 (full model) outperforms Group 4 in all three metrics. This indicates that Patch Token features enhance the model’s ability to understand semantic primitives. As noted in MaskCLIP [49], patch tokens retain local semantic details, which can help the model better understand fine-grained local attributes such as “dented” or “spiky”.
V-F Attention Analysis
Figure 9: Attention visualization. “CLS to Patch” visualizes the attention from the image’s CLS token to the patch tokens, highlighting salient visual regions. The subsequent columns (“Patch to Attribute”) visualize the attention from the image local region tokens to corresponding attribute text tokens.
In Fig.˜9, we conducted an attention visualization analysis to better understand the specific mechanism of the image patch tokens within our Integrator module. The attention maps of “Patch to Attribute” tend to exhibit a fixed pattern across different samples. They do not show a clear correspondence between each specific attribute and its associated image region. This is because patch tokens are not directly supervised for attribute alignment. This behavior is similar to that observed in GATv2[50] on fully connected graphs.
Despite the lack of explicit patch-level attribute alignment, the attention maps still reveal meaningful behavior of patch token. The “CLS to Patch” maps show the CLS token focusing on salient visual regions, such as the sliced bananas in the bag or the texture of the dried banana chips. This indicates that the main role of patch tokens is to provide rich local details to the global CLS token, which then performs the final prediction. While the model may not learn intuitive patch-level attribute alignment, it still effectively extracts and utilizes useful visual information from local regions.
TABLE VII: Impact of primitive interactions within the Integrator module. Results are obtained by retraining the model with specific attention interactions disabled: “Attr–Obj” denotes attention between attributes and objects, “Attr–Attr” denotes attention among attributes, and “All-Primitives” denotes attention over all primitive tokens.
Model Exact Match Attr Obj
Ours 17.42 71.55 86.45
Ours w/o Attr-Obj 15.25 69.04 86.87
Ours w/o Attr-Attr 16.97 70.67 86.68
Ours w/o All-Primitives 15.11 69.15 85.78
We also conducted an attention masking experiment to measure the impact of primitive relationships. As shown in Tab.˜VII, masking the interaction between attribute and object tokens (Ours w/o Attr-Obj) caused a significant drop in Exact Match and attribute accuracy. This drop was more pronounced than when we masked intra-attribute attention (Ours w/o Attr-Attr). This confirms that modeling the relationship between attributes and objects is more critical to our model’s performance than modeling the relationships among attributes themselves.
VI Conclusion